OpenStudy (anonymous):

Check my answer?? Two objects are launched horizontally from a cliff, one of them is launched at a much higher velocity. Which object will hit the ground first?

OpenStudy (anonymous):

The faster one The slower one THey will hit at the same time <--- My answer

OpenStudy (anonymous):

If they are both the same mass they would hit the ground at the same time. So yea your answer makes sense

OpenStudy (anonymous):

Yay! Thank you for answering! I got some more questions, do you mind helping me with some? @lilfrog2014

OpenStudy (anonymous):

I can try lol

OpenStudy (anonymous):

Yay! Thank you! What is the approximate instantaneous velocity of a freely falling object 10 seconds after it is released from rest, neglecting air resistance? (Round g to 10 m/s^2 instead of using the more precise value of 9.8 m/s^2)? 10 m/s 100 m/s <-- My answer 50 m/s 1000 m/s

OpenStudy (anonymous):

10m/s^2 is 10*10 so that's correct

OpenStudy (anonymous):

Yay awesome! Approximately how far will a freely falling object fall in the first 10 seconds after being released from rest, negelecting air resistance? (Approximate g as 10 m/s^2) 500m 100m <--- My answer 1000m 200m

OpenStudy (anonymous):

That's the same thing I think -- I'm gonna guess you'd be right given the information that they give in the question

OpenStudy (anonymous):

If anything you ask I end up getting wrong I'm sorry lol I'm not very good at this stuff but may I ask what grade you are in?

OpenStudy (anonymous):

@owlcatchyoulater24

OpenStudy (anonymous):

Im in 12th. I just dont understand the last two questions i asked because it seems like there's not enough information. Ill ask another question, and this time, do you think you could give me a formula to be able to solve it? Im just confused. Lol How tall is a cliff if a rock thrown horizontally from the top takes 4 seconds to hit the ground? 128.8 ft 257.6 ft 64.4 ft

OpenStudy (anonymous):

Umm hang on 1 sec lol

OpenStudy (anonymous):

I had a question on a quiz exactly like this once I wrote s= ut + 1/2 a t^2 s= 0 + 1/2 x 9.81 x 4^2 s= 78.48m which is 258.5 ft and that's closest to answer b

OpenStudy (anonymous):

Sorry had to find my old notes :p

OpenStudy (anonymous):

@owlcatchyoulater24 oh btw I'm only in 9th grade ... Lol

OpenStudy (anonymous):

Thank you! Oh man, you're smart! Lol!

OpenStudy (anonymous):

Lol no problem